OpenStudy (anonymous):

how do you find the derivative of h(x)=-x^3+15x^2-48x

OpenStudy (anonymous):

Apply (x^n)' = n x^ (n-1)

OpenStudy (campbell_st):

if y = x^n dy/dx = nx^(n-1) so h(x) = -x^3 + 15x^2 - 48x = \[h'(x) = -3\times x^{3-1} + 2\times15x^{2-1} - 4x^{1-1}\] gives \[h'(x) = -3x^2 + 30x - 48\]

OpenStudy (anonymous):

\[\LARGE h'(x)=(-x^3+15x^2-48x)'\] \[\LARGE h'(x)=-(x^3)'+(15x^2)'-(48x)'\] now apply \[\Huge (x^n)'=n\cdot x^{n-1 }\]

OpenStudy (radar):

Use the power rule. \[h ^{'}(x)=-3x ^{2}+30x-48\]